OpenStudy (anonymous):

Limit question here, I'm confused on how to get the answer:

OpenStudy (anonymous):

\[\lim \sqrt{9x ^{2}+ x} - 3x \] as x -> infinity

OpenStudy (anonymous):

I tried rationalizing the numerator and got: \[ \frac{ 9x ^{2} + 8x }{ \sqrt{9x ^{2}+x}+3x}\]

OpenStudy (anonymous):

and this is where i'm confused. I'm not quite sure what to do next. I know it has something to do with leading coefficients but i'm really confused.

OpenStudy (amistre64):

numerator?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Sorry, now i know what you meant. I tried putting the original problem into a fraction then rationalizing the numerator

OpenStudy (anonymous):

Was that a bad move?

OpenStudy (amistre64):

it might not be a bad move, but I cant really determine what the original problem was that needs its limit taken. you posted:\[\lim \sqrt{9x ^{2}+ x} - 3x\] this has no fraction in it to denote a numerator to rationalize.

OpenStudy (amistre64):

the conjugate of what you wrote is\[\sqrt{9x ^{2}+ x} + 3x\] and when multiplied we get\[{9x ^{2}+ x} - 9x^2=x\]